OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). A. y = -2x - 4 B. y = -2x - 8 C. y = -2x + 18 D. y = -2x – 12

hero (hero):

Substitute (x,y) =(2,-8) into the equations

hero (hero):

Appears to be A

OpenStudy (anonymous):

y + 8 = -2( x - 2) if memory serves me correct. that is the point slop form

OpenStudy (anonymous):

Thanks for that tip! :)

hero (hero):

You should state who you're directing the thanks towards

OpenStudy (anonymous):

Both of you! :)